# Solving rescheduling problems in heterogeneous urban railway networks using hybrid quantum-classical approach

### v3

@@ 4 Methods and Model
In the following we describe our model in detail. Section 4.1 defines sets, parameters and decision variables. Section 4.2 describes our Integer Linear Programming (ILP) formulation.
@@ 4.1 Sets, parameters and decisions
To formulate our decision variables, constraints and the objective function, we determine sets of index tuples needed to find the actual index sets of variables, from the given infrastructure data, timetable data, and the rolling stock circulation plan. Also, we introduce parameters calculated from the same input.
4.1.1 Sets
@@
Let us denote by $\hat{S}$ the set of all stations, and by $S\subset\hat{S}$ the decision stations we have determined in advance, as already described. The main objects of our model for railway rescheduling are the trains $j\in\mathcal{J}$ and the stations $s\in\mathcal{S}_{j}$ in their path. Set $\mathcal{S}_{j}$ includes decision stations only and it is an ordered set. In addition to $\mathcal{J}$ and $\mathcal{S}_{j}$, the relevant sets are the following. Set $\mathcal{J}^{2\,(\text{turn})}_{s}\subset\mathcal{J}\times\mathcal{J}\left(\forall s\in\mathcal{S}\right)$ is the set of all pairs of trains so that the first train of the pair terminates at station $s$ and its rolling stock continues as the second train of the pair. This set is deduced from the rolling stock circulation plan and the timetable. Set $\mathcal{J}^{2\,(\text{close})}\subset\mathcal{J}\times\mathcal{J}$ is the set of trains which are close enough to each other in time so that precedence variables have to be defined for them. More details will be given at the description of the parameter $d_{\max}$ which this set depends on. Set $\mathcal{J}^{2\,(\text{headway})}\subset\mathcal{J}^{2\,(\text{close})}$ is the set of train pairs that can potentially violate the minimal headway time between two subsequent trains on a line segment. For a pair of trains to be in $\mathcal{J}^{2\,(\text{headway})}$, both their routes need to include the same line segment so that the trains are moving in the same direction on it and can meet there according to model parameters, i.e. maximal allowed delay. Set $\mathcal{J}^{2\,(\text{single})}\subset\mathcal{J}^{2\,(\text{close})}$ is the set of all trains that share at least one single-track line segment as a common part of their routes so that they are heading in the opposite direction. Set $\left(\forall s\in\mathcal{S}\right)\ \mathcal{J}^{2\,(\text{track})}_{s}\subset\mathcal{J}^{2\,(\text{close})}$ is the set of train pairs that are planned to occupy the same track on station $s$ anytime during the planning horizon, thereby competing for the same track. Set $\left(\forall s\in\mathcal{S}\right)\ \mathcal{J}^{2\,(\text{switch},\text{out})}_{s}\subset\mathcal{J}^{2\,(\text{close})}$ is the set of train pairs that are planned to pass the same interlocking area of $s$ upon their departure from station $s$. Set $\left(\forall(s,s^{\prime})\in\mathcal{S}^{\times 2}\right)\ \mathcal{J}^{2\,(\text{switch},\text{out},\text{in})}_{s,s^{\prime}}\subset\mathcal{J}^{2\,(\text{close})}$ is the set of train pairs that are planned to pass the same interlocking area of $s$ in order to have $j$ to departure $s$ while $j^{\prime}$ arrive $s$ from the direction of $s^{\prime}$. Set $\left(\forall s\in\mathcal{S}^{\times 2}\right)\ \mathcal{J}^{2\,(\text{switch},\text{in},\text{noMP})}_{s,s^{\prime}}\subset\mathcal{J}^{2\,(\text{close})}$ is the set of train pairs that are planned to pass the same interlocking area of $s$ upon their arrival at $s$ from the direction of $s^{\prime}$, and there is no M-P possibility for them between $s$ and $s^{\prime}$ Set $\left(\forall s\in\mathcal{S}^{\times 2}\right)\ \mathcal{J}^{2\,(\text{switch},\text{in},\text{MP})}_{s,s^{\prime}}\subset\mathcal{J}^{2\,(\text{close})}$ is the set of train pairs that are planned to pass the same interlocking area of $s$ upon their arrival at $s$ so that either they both come from the direction of $s^{\prime}$ but there is a M-P possibility for them between $s$ and $s^{\prime}$, or one of them is approaching $s$ from a direction other than $s^{\prime}$. Set $\left(\forall j\in\mathcal{J}\right)\ \mathcal{C}^{2}_{j}$ is the set of all station pairs that are subsequent in the route of $j$. Set $\left(\forall(j,j^{\prime})\in\mathcal{J}^{\times 2}\right)\ \mathcal{C}^{2\,(\text{common})}_{j,j^{\prime}}$ is the set of all station pairs that appear as subsequent stations in the route of both $j$ and $j^{\prime}$ heading in the same direction. Set $\left(\forall(j,j^{\prime})\in\mathcal{J}^{\times 2}\right)\ \mathcal{C}^{2\,(\text{common, single})}_{j,j^{\prime}}$ is the set of all station pairs that appear as subsequent stations in the route of both $j$ and $j^{\prime}$ which are connected with a single-track line segment, and trains are heading in opposite directions. The order within the pairs is determined by $j$. All these sets can be enumerated on the basis of the input data in a straightforward manner.
4.1.2 Parameters
@@
  • $\upsilon(j,s)$ is the earliest possible departure time of $j$ from $s$ when a disturbance and/or disruption in the network happens. It is the maximum of $\sigma(j,s)$ and the technically feasible earliest departure time, and may be more constraining than the first depending on the initial delays. The latter is calculated assuming that the train follows its planned route at minimum running time, and that there are no other trains present on the network.
  • $d_{\text{max}}$ is an upper bound assumed for the secondary delay $t^{(\text{out})}(j,s)-\upsilon(j,s)$. This parameter sets an upper limit for the possible secondary delays that can arise on the particular part of the network. Setting such a bound is common in the literature, see e.g. the work of [14]. We set the same value of this parameter for all our testing instances. It has to be big enough so that no secondary delay exceeds it (e.g. $d_{\max}=40$ excludes the possibility of a one-hour delay due to waiting for another delayed train). Meanwhile it is desirable to set its value as small as possible; restricting the time variables to small intervals decreases the number of binary decision variables in the model.
The choice of a small $d_{\max}$ parameter results in a decrease of the model size, which is achieved via defining the set $J^{2\,(\text{close})}$. A pair of trains $(j,j^{\prime})$ is included in this set if and only if they can meet on any station, given the original timetable, the disturbed/disrupted timetable, and assuming that no train can have a secondary delay greater than $d_{\max}$. The number of constraints will be proportional to the average number of trains that can meet another train at a station; this will be linear in $d_{\max}$.
4.1.3 Decision variables
@@
Second, in addition to the time variables, we use three sets of binary precedence variables. Decision variables $y^{(\text{out})}(j,j^{\prime},s)\in\{0,1\}$ determine the order of trains to leave stations: the variable takes the value $1$ if train $j$ leaves station $s$ before train $j^{\prime}$, and 0 otherwise. The $(j,j^{\prime},s)$ tuples for which an $y$ variable is defined will be specified later. Similarly, the precedence variables $y^{(\text{in})}(j,j^{\prime},s)\in\{0,1\}$ prescribe the order at the entry to stations; the value is $1$ if $j$ arrives to $s$ before $j^{\prime}$. The last set of precedence variables describes the precedence of trains at some resource located between stations $s$ and $s^{\prime}$ (e.g. single track used by $j$ and $j^{\prime}$ heading in opposite direction). The binary variable $z(j,j^{\prime},s,s^{\prime})\in\{0,1\}$ will be $1$ if $j$ uses the given resource before $j^{\prime}$. Also in this case, the quadruples $(j,j^{\prime},s,s^{\prime})$ for which we have such a variable will be specified later.
@@ 4.2.2 Constraints
where $C$ is a constant big enough to make the constraint satisfied whenever the binary variable $y^{(\text{out})}(j^{\prime},j,s)$ takes the value of $1$. (Constraints of this structure are often termed as big-M constraints in this context.) In our implementation we calculate and use the smallest suitable value of $C$ given particular $d_{\max}$ value, i.e.
@@
$\small{\forall s\in S\forall(j,j^{\prime})\in\mathcal{J}^{2\,(\text{track})}_{s}}t^{(\text{in})}(j^{\prime},s)\geq t^{(\text{out})}(j,s)-C\cdot y^{(\text{out})}(j^{\prime},j,s),\$ (9)
@@
These are introduced in order to bind the train with the shunting movement called also service train. If the train set of train $j$ which terminates at $s$ is supposed to continue its trip as (service train) $j^{\prime}$ or vice versa, a precedence of these trains including a minimum turnaround time has to be ensured:
@@
As for the scaling of our model, the number of time variables i.e. #$\#t$ is bounded by number of trains times number of stations #$\#\mathcal{J}\#\mathcal{S}$ (see also Section 3.1 of the work of  [21]). In order to obtain a more accurate estimate of the number of these variables we observe that in our model no trains visit all stations. We assume that in average each train visits $\alpha\in(0,1)$ fraction of the stations, we have then
@@
Suppose, that most of the network is composed mainly of double track lines. Then there will be typically a single precedence variable per train and station ($y^{(out)}(j,j^{\prime},s)$), and thus their number can be estimated by:
@@
where $N(d_{\max})$ is number of trains the train can meet at a station on average. Under such assumptions the total number of variables grows linearly with number of trains and stations:
@@
If, on the other hand, the network has dominantly single track lines, two precedence variables per train and station ($y^{(out)}(j,j^{\prime},s)$ and $z(j,j^{\prime},s,s^{\prime})$) will be needed, and thus
@@
Based on Eq. (16) there are typically $2$ constraints per each $y$ variable as headway and station track occupation constraints, and $2$ constraints per each $z$ variable as single track occupation constraints. For the interlocking area occupation constraints it is more complicated as it involves both $z$ and $y$ variables; a good assumption is to consider $2$ constraints for each $y$ variable. For the running time and minimal dwell time we expect one constraint per train and station. Thus number of constraints can be estimated by
@@
where the second term is tied to single track line conditions (one condition per $z$ variable).
#$\#constr.\approx\begin{cases}\alpha(6N(d_{\max})+2)\#\mathcal{J}\#\mathcal{S}\text{ for double track}\\ \alpha(8N(d_{\max})+2)\#\mathcal{J}\#\mathcal{S}\text{ for single track}.\end{cases}$ (22)
Because of the limitations in time imposed by $d_{\max}$ the model can be viewed as linear in number of trains and number of stations. Hence, as expected, the model is more complex for the network in which single track lines dominate. Although the analysed network is mostly composed of the double track lines, we will analyse certain use-cases with the fraction of single track lines increased.
Observe that without limiting the secondary delays by $d_{\max}$, the value of $N(d_{\max})$ would only be bounded by #$\#\mathcal{J}-1$ (as each train is considered to possibly meet each other), which would result in quadratic scaling of the number of precedence variables and number of constraints in the number of trains.
@@ 5 Hybrid quantum-based approach
To overcome the limitations described in in Section II of current hardware quantum annealers, we apply hybrid (quantum classical) solver. In particular, we use the ’Leap Hybrid Solver Service (HSS)’ [11]; a cloud-based proprietary solver which is developed by the market leader of QA hardware and is available as a service. It utilizes a hybrid approach that combines classical computational power with quantum processing.
In particular, we have used the Constrained Quadratic Model (CQM) [12] Solver. The CQM hybrid solver works as follows.
@@
A constrained linear or quadratic model, and the value of the solver parameter t_min, which prescribes the minimum required run time (in seconds) solver is allowed to work on any problem.
@@
Preprocessing includes subproblem identification and decomposition to smaller instances [52]. Furthermore, the solver’s internal preprocessing mechanism takes constraints into account automatically, implementing the suitable penalty method internally.
@@
The hybrid solver implements a workflow that combines a portfolio of various classical heuristics (including tabu search, simulated annealing, etc.) running on powerful CPUs (Central Processing Units) and GPUs (Graphics Processing Units). In the course of the solution process the hardware quantum annealer is also invoked, in order to solve smaller QUBO problems that can potentially boost the classical heuristics. Integer problems are transformed to binary using state-of-the-art methods like the one described by [25][34], generally applicable for integer problems. When using the quantum hardware, the solver performs the required embedding, runs the QUBO subproblem on the physical QA device several times, and returns a sample of potential solutions. After these readouts, the sample is incorporated into the solution workflow. In this way, even though the physical annealer supports problems with limited size (e.g. the Pegasus machine has $5500$ physical qubits [16], and several of these may be needed to represent a binary variable because of the embedding), the approximate solutions of small subproblems contribute to the solution process. The final solution is composed by the solver’s module.
@@
The solution of the integer or quadratic constrained problem. The workflow results in a solution that is not guaranteed to be optimal, hence, the hybrid solver as a whole is in nature a heuristic solver. Furthermore, the solver is probabilistic, hence each at each run the final solution may be different.
While D-Wave’s proprietary solvers act as black boxes, i.e. they hide the exact details of the described process process from the user, the output is supplemented with timing parameters, including run time: the total elapsed time including system overhead, and Quantum Processing Unit (QPU) access time which is is the time spent accessing the actual quantum hardware. These parameters enable a comparison with other solvers.
As mentioned before, in addition to the actual problem to be solved, the CQM solver optionally inputs another parameter, t_min. This is a time limit for heuristics running in parallel[12]: unless a thread does not terminate by itself after t_min, it is allowed to run at least till t_min before it is stopped at the current best solution. We have sampled the CQM problems with various settings of t_min and uncovered the impact of this setting to the model performance and solution quality for our problem.
The CQM Solver can solve problems encoded in the form of Eq. (2) – (15) (c.f. Section 4.2) with up to $5000$ binary or integer variables and $100,000$ constraints. Computational results of railway rescheduling problems were obtained using classical as well as hybrid (quantum classical) solvers. As a classical solver, IBM ILOG CPLEX Interactive Optimizer (version 22.1.0.0) was used. The CPLEX computation was performed on $16$ cores of Intel(R) Core(TM) i7-10700kF CPU 3.80GHz with $64$GBs of memory. As hybrid solver, the D-Wave CQM hybrid solver (version: 1.12) was used.

### v4

@@ 4 Methods and Model
In the following, we describe our model in detail. Section 4.1 defines sets, parameters, and decision variables. Section 4.2 describes our Integer Linear Programming (ILP) formulation.
@@ 4.1 Sets, parameters and decisions
To formulate our decision variables, constraints, and the objective function, we determine sets of index tuples needed to find the actual index sets of variables, from the given infrastructure data, timetable data, and the rolling stock circulation plan. Also, we introduce parameters calculated from the same input.
Sets
@@
Let us denote by $\hat{S}$ the set of all stations, and by $S\subset\hat{S}$ the decision stations we have determined in advance, as already described. The main objects of our model for railway rescheduling are the trains $j\in\mathcal{J}$ and the stations $s\in\mathcal{S}_{j}$ in their path. Set $\mathcal{S}_{j}$ includes decision stations only and it is an ordered set. In addition to $\mathcal{J}$ and $\mathcal{S}_{j}$, the relevant sets are the following. Set $\mathcal{J}^{2\,(\text{turn})}_{s}\subset\mathcal{J}\times\mathcal{J}\left(\forall s\in\mathcal{S}\right)$ is the set of all pairs of trains so that the first train of the pair terminates at station $s$ and its rolling stock continues as the second train of the pair. This set is deduced from the rolling stock circulation plan and the timetable. Set $\mathcal{J}^{2\,(\text{close})}\subset\mathcal{J}\times\mathcal{J}$ is the set of trains that are close enough to each other in time so that precedence variables have to be defined for them. More details will be given in the description of the parameter $d_{\max}$ which this set depends on. Set $\mathcal{J}^{2\,(\text{headway})}\subset\mathcal{J}^{2\,(\text{close})}$ is the set of train pairs that can potentially violate the minimal headway time between two subsequent trains on a line segment. For a pair of trains to be in $\mathcal{J}^{2\,(\text{headway})}$, both their routes need to include the same line segment so that the trains are moving in the same direction on it and can meet there according to model parameters, i.e. maximal allowed delay. Set $\mathcal{J}^{2\,(\text{single})}\subset\mathcal{J}^{2\,(\text{close})}$ is the set of all trains that share at least one single-track line segment as a common part of their routes so that they are heading in the opposite direction. Set $\left(\forall s\in\mathcal{S}\right)\ \mathcal{J}^{2\,(\text{track})}_{s}\subset\mathcal{J}^{2\,(\text{close})}$ is the set of train pairs that are planned to occupy the same track on station $s$ anytime during the planning horizon, thereby competing for the same track. Set $\left(\forall s\in\mathcal{S}\right)\ \mathcal{J}^{2\,(\text{switch},\text{out})}_{s}\subset\mathcal{J}^{2\,(\text{close})}$ is the set of train pairs that are planned to pass the same interlocking area of $s$ upon their departure from station $s$. Set $\left(\forall(s,s^{\prime})\in\mathcal{S}^{\times 2}\right)\ \mathcal{J}^{2\,(\text{switch},\text{out},\text{in})}_{s,s^{\prime}}\subset\mathcal{J}^{2\,(\text{close})}$ is the set of train pairs that are planned to pass the same interlocking area of $s$ in order to have $j$ to departure $s$ while $j^{\prime}$ arrive $s$ from the direction of $s^{\prime}$. Set $\left(\forall s\in\mathcal{S}^{\times 2}\right)\ \mathcal{J}^{2\,(\text{switch},\text{in},\text{noMP})}_{s,s^{\prime}}\subset\mathcal{J}^{2\,(\text{close})}$ is the set of train pairs that are planned to pass the same interlocking area of $s$ upon their arrival at $s$ from the direction of $s^{\prime}$, and there is no M-P possibility for them between $s$ and $s^{\prime}$ Set $\left(\forall s\in\mathcal{S}^{\times 2}\right)\ \mathcal{J}^{2\,(\text{switch},\text{in},\text{MP})}_{s,s^{\prime}}\subset\mathcal{J}^{2\,(\text{close})}$ is the set of train pairs that are planned to pass the same interlocking area of $s$ upon their arrival at $s$ so that either they both come from the direction of $s^{\prime}$ but there is a M-P possibility for them between $s$ and $s^{\prime}$, or one of them is approaching $s$ from a direction other than $s^{\prime}$. Set $\left(\forall j\in\mathcal{J}\right)\ \mathcal{C}^{2}_{j}$ is the set of all station pairs that are subsequent in the route of $j$. Set $\left(\forall(j,j^{\prime})\in\mathcal{J}^{\times 2}\right)\ \mathcal{C}^{2\,(\text{common})}_{j,j^{\prime}}$ is the set of all station pairs that appear as subsequent stations in the route of both $j$ and $j^{\prime}$ heading in the same direction. Set $\left(\forall(j,j^{\prime})\in\mathcal{J}^{\times 2}\right)\ \mathcal{C}^{2\,(\text{common, single})}_{j,j^{\prime}}$ is the set of all station pairs that appear as subsequent stations in the route of both $j$ and $j^{\prime}$ which are connected with a single-track line segment, and trains are heading in opposite directions. The order within the pairs is determined by $j$. All these sets can be enumerated based on the input data in a straightforward manner.
Parameters
@@
  • $\upsilon(j,s)$ is the earliest possible departure time of $j$ from $s$ when a disturbance in the network happens. It is the maximum of $\sigma(j,s)$ and the technically feasible earliest departure time, and may be more constraining than the first depending on the initial delays. The latter is calculated assuming that the train follows its planned route at minimum running time and that there are no other trains present on the network.
  • $d_{\text{max}}$ is an upper bound assumed for the secondary delay $t^{(\text{out})}(j,s)-\upsilon(j,s)$. This parameter sets an upper limit for the possible secondary delays that can arise on a particular part of the network. Setting such a bound is common in the literature, see e.g. the work of D’Ariano et al. (2007). We set the same value of this parameter for all our testing instances. It has to be big enough so that no secondary delay exceeds it (e.g. $d_{\max}=40$ excludes the possibility of a one-hour delay due to waiting for another delayed train). Meanwhile, it is desirable to set its value as small as possible; restricting the time variables to small intervals decreases the number of binary decision variables in the model.
The choice of a small $d_{\max}$ parameter results in a decrease of the model size, which is achieved via defining the set $J^{2\,(\text{close})}$. A pair of trains $(j,j^{\prime})$ is included in this set if and only if they can meet on any station, given the original timetable, the disturbed timetable, and assuming that no train can have a secondary delay greater than $d_{\max}$. The number of constraints will be proportional to the average number of trains that can meet another train at a station; this will be linear in $d_{\max}$.
Decision variables
@@
Second, in addition to the time variables, we use three sets of binary precedence variables. Decision variables $y^{(\text{out})}(j,j^{\prime},s)\in\{0,1\}$ determine the order of trains to leave stations: the variable takes the value $1$ if train $j$ leaves station $s$ before train $j^{\prime}$, and 0 otherwise. The $(j,j^{\prime},s)$ tuples for which an $y$ variable is defined will be specified later. Similarly, the precedence variables $y^{(\text{in})}(j,j^{\prime},s)\in\{0,1\}$ prescribe the order at the entry to stations; the value is $1$ if $j$ arrives to $s$ before $j^{\prime}$. The last set of precedence variables describes the precedence of trains at some resource located between stations $s$ and $s^{\prime}$ (e.g. a single-track segment used by $j$ and $j^{\prime}$ heading in opposite direction). The binary variable $z(j,j^{\prime},s,s^{\prime})\in\{0,1\}$ will be $1$ if $j$ uses the given resource before $j^{\prime}$. Also in this case, the quadruples $(j,j^{\prime},s,s^{\prime})$ for which we have such a variable will be specified later.
@@ 4.2 ILP formulation
where $C$ is a constant big enough to make the constraint satisfied whenever the binary variable $y^{(\text{out})}(j^{\prime},j,s)$ takes the value of $1$. (Constraints of this structure are often termed as big-M constraints in this context.) In our implementation, we calculate and use the smallest suitable value of $C$ given particular $d_{\max}$ value, i.e.
@@
${\forall s\in S\forall(j,j^{\prime})\in\mathcal{J}^{2\,(\text{track})}_{s}}t^{(\text{in})}(j^{\prime},s)\geq t^{(\text{out})}(j,s)-C\cdot y^{(\text{out})}(j^{\prime},j,s),\$ (9)
@@
These are introduced to bind the train with the shunting movement called also service train. If the train set of train $j$ which terminates at $s$ is supposed to continue its trip as (service train) $j^{\prime}$ or vice versa, a precedence of these trains including a minimum turnaround time has to be ensured:
@@
As for the scaling of our model, the number of time variables i.e. #$\#t$ is bounded by number of trains times number of stations #$\#\mathcal{J}\#\mathcal{S}$ (see also Section 3.1 of the work of  Domino et al. (2022b)). To obtain a more accurate estimate of the number of these variables we observe that in our model no trains visit all stations. We assume that in average each train visits $\alpha\in(0,1)$ fraction of the stations, we have then
@@
Suppose, that most of the network is composed mainly of double-track lines. Then there will be typically a single precedence variable per train and station ($y^{(out)}(j,j^{\prime},s)$), and thus their number can be estimated by:
@@
where $N(d_{\max})$ is the number of trains the train can meet at a station on average. Under such assumptions, the total number of variables grows linearly with number of trains and stations:
@@
If, on the other hand, the network has dominantly single-track lines, two precedence variables per train and station ($y^{(out)}(j,j^{\prime},s)$ and $z(j,j^{\prime},s,s^{\prime})$) will be needed, and thus
@@
Based on Eq. (16) there are typically $2$ constraints per each $y$ variable as headway and station track occupation constraints, and $2$ constraints per each $z$ variable as single-track occupation constraints. For the interlocking area occupation constraints, it is more complicated as it involves both $z$ and $y$ variables; a good assumption is to consider $2$ constraints for each $y$ variable. For the running time and minimal dwell time we expect one constraint per train and station. This number of constraints can be estimated by
@@
where the second term is tied to single-track line conditions (one condition per $z$ variable).
#$\#constr.\approx\begin{cases}\alpha(6N(d_{\max})+2)\#\mathcal{J}\#\mathcal{S}\text{ for double-track}\\ \alpha(8N(d_{\max})+2)\#\mathcal{J}\#\mathcal{S}\text{ for single-track}.\end{cases}$ (22)
Because of the limitations in time imposed by $d_{\max}$ the model can be viewed as linear in number of trains and number of stations. Hence, as expected, the model is more complex for the network in which single-track lines dominate. Although the analyzed network is mostly composed of the double-track lines, we will analyze certain use-cases with the fraction of single-track lines increased.
Observe that without limiting the secondary delays by $d_{\max}$, the value of $N(d_{\max})$ would only be bounded by #$\#\mathcal{J}-1$ (as each train is considered to possibly meet each other), which would result in quadratic scaling of the number of precedence variables, and that of the number of constraints in the number of trains.
@@ 5 Hybrid quantum-based approach
To overcome the limitations described in in Section 2 of current hardware quantum annealers, we apply a hybrid (quantum-classical) solver. In particular, we use the ’Leap Hybrid Solver Service (HSS)’ (D-Wave Quantum Inc., 2022a); a cloud-based proprietary solver that is developed by the market leader of QA hardware and is available as a service. It utilizes a hybrid approach that combines classical computational power with quantum processing.
In particular, we have used the Constrained Quadratic Model (CQM) (D-Wave Quantum Inc., 2022b) Solver. The CQM hybrid solver works as follows.
@@
A constrained linear or quadratic model, and the value of the solver parameter t_min, which prescribes the minimum required run time (in seconds) the solver is allowed to work on any problem.
@@
Preprocessing includes subproblem identification and decomposition to smaller instances (Tran et al., 2016). Furthermore, the solver’s internal preprocessing mechanism takes constraints into account automatically, implementing the suitable penalty method internally.
@@
The hybrid solver implements a workflow that combines a portfolio of various classical heuristics (including tabu search, simulated annealing, etc.) running on powerful CPUs (Central Processing Units) and GPUs (Graphics Processing Units). In the course of the solution process the hardware quantum annealer is also invoked, to solve smaller QUBO problems that can potentially boost the classical heuristics. Integer problems are transformed to binary using state-of-the-art methods like the one described by Glover et al. (2019); Karimi and Ronagh (2019), generally applicable for integer problems. When using the quantum hardware, the solver performs the required embedding, runs the QUBO subproblem on the physical QA device several times, and returns a sample of potential solutions. After these readouts, the sample is incorporated into the solution workflow. In this way, even though the physical annealer supports problems with limited size (e.g. the Pegasus machine has $5500$ physical qubits (Dattani et al., 2019), and several of these may be needed to represent a binary variable because of the embedding), the approximate solutions of small subproblems contribute to the solution process. The final solution is composed by the solver’s module.
@@
The solution or solutions of quadratic constrained problem or the integer problem. The workflow results in a solution or solutions that are not guaranteed to be optimal, hence, the hybrid solver as a whole is in nature a heuristic solver. Furthermore, the solver is probabilistic, hence at each run the final output can be different. The CQM (integer) solver returns a series of solutions, which can be an additional asset. Nevertheless, for sake of presentation, our approach is to analyze only the best CQM solution given each run of the hybrid algorithm (realization).
While D-Wave’s proprietary solvers act as black boxes, i.e. they hide the exact details of the described process from the user, the output is supplemented with timing parameters, including run time: the total elapsed time including system overhead, and Quantum Processing Unit (QPU) access time which is the time spent accessing the actual quantum hardware. These parameters enable a comparison with other solvers.
As mentioned before, in addition to the actual problem to be solved, the CQM solver optionally inputs another parameter, t_min. This is a time limit for heuristics running in parallel(D-Wave Quantum Inc., 2022b): unless a thread does not terminate by itself after t_min, it is allowed to run at least till t_min before it is stopped at the current best solution. We have sampled the CQM problems with various settings of t_min and uncovered the impact of this setting on the model performance and solution quality for our problem.
The CQM Solver can solve problems encoded in the form of Eq. (2) – (15) (c.f. Section 4.2) with up to $5000$ binary or integer variables and $100,000$ constraints. Computational results of railway rescheduling problems were obtained using classical as well as hybrid (quantum-classical) solvers. As a classical solver, IBM ILOG CPLEX Interactive Optimizer (version 22.1.0.0) was used. The CPLEX computation was performed on $16$ cores of Intel(R) Core(TM) i7-10700kF CPU 3.80GHz with $64$GBs of memory. As a hybrid solver, the D-Wave CQM hybrid solver (version: 1.12) was used.